Hanna, who is a 5-year-old girl, eats nothing but pasta, yogurt, and lemonade. Each month her parents buy 25 pounds of pasta, 65 packages of yogurt, and 15 bottles of lemonade. These three items make up a basket of goods and services similar to the much larger basket used by the Bureau of Labor Statistics (BLS) when computing the official Consumer Price Index (CPI). The table below lists the average price for each item in this basket for the past four years.

Hanna's Meals

Year Pasta (dollars per pound) Yogurt (dollars per package) Lemonade (dollars per bottle)
1 $1.90 $1.00 $2.10
2 2.10 1.10 2.20
3 2.25 1.10 1.95
4 2.20 1.20 2.00

a. For each year, calculate the CPI, using year 1 as the base year.

In year 1, the CPI was  .

In year 2, the CPI was  .

In year 3, the CPI was  .

In year 4, the CPI was  .

b. For each year, calculate the CPI, using year 3 as the base year.

In year 1, the CPI was  .

In year 2, the CPI was  .

In year 3, the CPI was  .

In year 4, the CPI was  .
Answer #1

Solution:-

First calculate the cost of basket in each year,

Year

Price of Pasta

Price of Yogurt

Price of Lemonade

Cost of basket

1

1.9

1

2.1

(25 *1.9 + 65 * 1 + 15 * 2.1) = 144

2

2.1

1.1

2.2

(25*2.1 + 65 * 1.1 + 15 * 2.2) = 157

3

2.25

1.1

1.95

(25 * 2.25 + 65 * 1.1 + 15 * 1.95) = 157

4

2.2

1.2

2

(25 * 2.2 + 65 * 1.2 + 15 * 2) = 163

Cost of Basket = (25 * Price of Pasta in year n) + (65 * Price of Yogurt in year n) + (15 * Price of Lemonade in year n)

Now calculate the CPI values using the following formula:-

CPI in Year n = Cost of the basket in year n / Cost of the basket in base year * 100

Cost Of Basket

CPI using year 1 as a base year

CPI using year 3 as a base year

144

(144 / 144 * 100) = 100

(144 / 157 * 100) = 91.72

157

(157 / 144 * 100) = 109.03

(157 / 157 * 100) = 100

157

(157 / 144 * 100) = 109.03

(157 / 157 * 100) = 100

163

(163 / 144 * 100) = 113.19

(163 / 157 * 100) = 103.82

a. For each year, calculate the CPI, using year 1 as the base year.

In year 1, the CPI was = 100

In year 2, the CPI was = 109.3

In year 3, the CPI was = 109.3

In year 4, the CPI was = 113.19

b. For each year, calculate the CPI, using year 3 as the base year.

In year 1, the CPI was = 91.72

In year 2, the CPI was = 100

In year 3, the CPI was = 100

In year 4, the CPI was = 103.82